OpenStudy (anonymous):

How do you write the standard number -759.82 in exponential notation?

OpenStudy (anonymous):

-7(102) + -5(101) + -9(100) + -8(10-1) + -2(10-2) -7(104) + -5(103) + -9(102) + -8(101) + -2(100) -7(103) + -5(102) + -9(101) + -8(100) + -2(10-1) -7(101) + -5(100) + -9(10-1) + -8(10-2) + -2(10-3)

OpenStudy (michele_laino):

hint: \[759.82 = 700 + 50 + 9 + 0.8 + 0.02\]

OpenStudy (anonymous):

-7(102) + -5(101) + -9(100) + -8(10-1) + -2(10-2) this one?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thanks, do you think you could help me with more?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Write the number 3,430 x 10-8 in scientific notation.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: \[3,430 = 3.4 \times {10^3}\]

OpenStudy (michele_laino):

oops...I have made an error:

OpenStudy (michele_laino):

\[3,430 = 3.43 \times {10^3}\]

OpenStudy (anonymous):

34.3 x 10-5

OpenStudy (michele_laino):

3.43 x 10^(-5)

OpenStudy (michele_laino):

since we have: \[3,430 \times {10^{ - 8}} = 3.43 \times {10^3} \times {10^{ - 8}} = 3.43 \times {10^{ - 5}}\]

OpenStudy (anonymous):

Oh my bad, put the decimal in the wrong place!